OpenStudy (anonymous):

Use S = n2 to find the sum of 1 + 3 + 5 + . . . + 499

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

a(n)=1+2(n-1) 499=1+2(k-1) 498=2(k-1) 249=k-1 k=248

OpenStudy (aum):

S = 1 + 3 + 5 + . . . + 499 This is an arithmetic series with first term \(a_1 = 1\), common difference \(d = 2\) and the last term 499. First, find how many terms there are in the series. The nth term of an arithmetic sequence is given by the formula: \(a_n = a_1 + (n-1)d\). Therefore, \[499 = 1 + (n-1)*2 \\ 498 = 2(n-1) \\ 249 = n - 1 \\ n = 250 \]The sum of the series is given by the formula: \(S_n = (a_1 + a_{last})/2 * n = (1 + 499)/2 * 250 = ~?\)